OpenStudy (anonymous):

I have one question that I really need help with, please help me!!! If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the first equation. x – 6y + 3z = 0 2x – 5y + 8z = 7 x + 7y – 4z = 13

OpenStudy (anonymous):

Solve the last two equations for y and z to obtain \[ \begin{array}{cc} y=\frac{1}{9} (33-4 x) &\\ z=\frac{19}{6}-\frac{19 x}{36} \end{array} \]

OpenStudy (anonymous):

Replace y and z by their values in equation 1 to obtain \[-\frac{2}{3} (33-4 x)+3 \left(\frac{19}{6}-\frac{19 x}{36}\right)+x=0\\ \] After simplification, you get \[ x=6 \]

OpenStudy (anonymous):

The new set of equations will be \[ x=6\\ \begin{array}{cc} y=\frac{1}{9} (33-4 x) &\\ z=\frac{19}{6}-\frac{19 x}{36} \end{array} \]